Shift report 20030207 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Feb 07 2003 - 15:52:49 EST

  • Next message: Flemming Videbaek: "RHIC running"
    ------------------------------------------------------------
    Date: Fri Feb  7 15:52:49 2003, Shift: 16:00-24:00
    Supervisor: PS, with: MG
    Log book#14 pages 4 - 4
    
    Shift Summary:
    ------------------------------------------------------------
    
    1. We had one long store during the shift. The store
    started around 9:20 and lasted until 13:00. The beam was restored around 14:15 but was dumped after a short time.
    
    2. Runs 7820, 7822, 7822, 7823, 7824 are data runs
    with MRS at 35 deg (A500) and with FS at 8 deg. (1/15B).
    These runs contain 500k MRS triggers and 98k FFS triggers.
    run 7626 is none TPC pedestal run
    run 7627 in TMC pedestal run, and
    run 7628 in DVM run.
    
    3. T3 and T5 tripped ones during a high background. 
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Fri Feb 07 2003 - 16:07:13 EST